# Group Therapy for Halos: Advancing Halo Mass Estimation for Galaxy Groups

### v1

@@ 4 Observational Application of Halo Masses
Figure 5: Empirical halo mass function (HMF) for galaxy groups with three or more members, constructed from the group catalogue of Van Kempen et al. (2024). The HMF is shown separately for groups in the SGP region dominated by 2dF coverage (red triangles), the G23 region with GAMA spectroscopy (orange diamonds), and the G23 region with 2dF spectroscopy (green crosses). Halo masses are estimated using the traditional and calibrated virial theorem method ($\log M_{\mathrm{halo,\,VT}}$ or $\log M_{\mathrm{halo,\,CVT}}$). Number densities are not corrected for survey volume or selection effects, due to the heterogeneous nature of the group sample and the challenges in defining a complete selection function. The solid black curve represents the analytic HMF prediction for the adopted cosmology of Shark, while the grey dashed line shows the empirical fit from Driver et al. (2022a) based on GAMA5, SDSS5, and REFLEX II data. Error bars reflect Poisson uncertainties.
The calibrated halo mass estimators developed in this work enable a range of new applications in the analysis of spectroscopic group catalogues. Here, we present two representative examples using the SGP group sample from Van Kempen et al. (2024): (i) the construction of an empirical HMF using the calibrated virial theorem, and (ii) the mapping of the quenched fraction of galaxies in the stellar mass–halo mass plane using the SHMR-based halo mass estimates.
@@ 4.1 Halo Mass Function from Observational Groups
The HMF is a fundamental cosmological observable, that encodes information about the growth of structure and the underlying cosmological parameters (e.g., Press & Schechter, 1974; Jenkins et al., 2001; Murray et al., 2013). In Figure 5, we present the empirical HMF for galaxy groups with $N_{\mathrm{m}}\geq 4$ members. The sample is divided into three subsamples: the SGP - 2dF (red triangles) which utilises 2dFGRS photometry across the entire sample, the G23 - 2dF region (green crosses) which utilises 2dFGRS photometry in the G23 region, and the G23 - GAMA (orange diamonds) which utilises the GAMA photometry.
@@
Unlike the SMF, which is based on uniform measurements of individual galaxies, constructing the HMF requires consistent data for all members within each group and a well-defined selection function to enable $1/V_{\mathrm{max}}$ corrections. However, the observational group sample is heterogeneous, with group members often drawn from multiple surveys that have varying completeness and magnitude limits (see Van Kempen et al., 2024). This diversity makes it challenging to accurately determine the completeness for each group, and thus we do not apply the $1/V_{\mathrm{max}}$ correction to the observational HMF. The absence of this correction means that the HMF is not fully corrected for incompleteness, particularly at the low-mass end. Low-mass groups are more likely to be missed because their member galaxies often fall below the survey’s limiting magnitude, and even when detected, such groups are typically found only at very low redshifts, where the survey volume is small and statistical uncertainties are larger. As a result, the observed HMF underestimates the true abundance of low-mass haloes, leading to an artificial drop-off in this regime. Therefore, the HMF presented here should be regarded as a demonstration of the methodology rather than a definitive measurement.
To ensure reliable binning of the HMF, the provided HMFs includes groups with $N_{\mathrm{m}}\geq 3$ detected in each survey (SGP-2dF, G23-2dF, and G23-GAMA). For the SGP-2dF sample, an additional restriction to $z<0.08$ is imposed to minimise the effects of declining completeness at higher redshifts.
@@
The HMFs shown in Figure 5 are derived using both the standard (left panel: Equation 3) and calibrated virial theorem (right panel: Equation 6) estimators. As discussed in Section 3.1.1 and illustrated in Figure 3, the standard virial theorem estimator predicts an excess of low-mass haloes. When considering the observational survey volume and the presence of a local void at $z\sim 0.02$ (Driver et al., 2022b; Van Kempen et al., 2024), such low-mass systems are not expected to be sampled within this dataset. This underestimation of halo masses by the traditional virial theorem could be misinterpreted as a genuine feature in the absence of this prior knowledge. In contrast, the calibrated virial theorem corrects for this bias, yielding a halo mass function that is more consistent with theoretical and practical expectations, providing a more reliable representation of the underlying halo population.
@@
The resulting HMFs are compared both to the analytic prediction for the adopted cosmology of Shark and to the empirical fit derived by Driver et al. (2022a). It is important to note that the dataset utilised by Driver et al. (2022a) to establish their empirical relation is substantially larger and more homogeneous than the sample analysed here. Nevertheless, our non-idealised and considerably smaller dataset yields a HMF that aligns well with theoretical expectations at the high to moderate halo mass range ($M_{halo}>12.75$) end underscores the significant progress made in the development of accurate halo mass estimators. However, at lower halo masses, the observed HMFs exhibit clear signs of incompleteness, further emphasising the necessity for larger and more homogeneous samples to robustly characterise this regime and improve observational constraints on the low-mass end.
Looking ahead, upcoming wide-area spectroscopic surveys such as WAVES (Driver et al., 2019) and 4HS (Taylor et al., 2023) will provide the depth, completeness, and sky coverage required to construct high-precision HMFs over a wide mass range. With sufficiently large samples and accurate halo mass estimates, such surveys will enable the use of the HMF as a cosmological probe, constraining parameters such as $\Omega_{m}$, $\sigma_{8}$, and potentially the nature of dark energy (e.g., Murray et al., 2013; Bocquet et al., 2019). The principal limiting factor for future HMF-based cosmological analyses is likely to be the uncertainty in group identification and membership assignment, particularly at low multiplicity, rather than the precision of the halo mass estimator itself.
@@ 4.2 Quenched Fraction in the Stellar Mass–Halo Mass Plane
Figure 7: Distribution of quenched fraction in the stellar mass–halo mass plane for galaxy groups in the observational dataset of Van Kempen et al. (2024). The colour scale represents the fraction of quenched galaxies (defined as those with $\log\,\mathrm{sSFR}<-11.0$) within each region of parameter space, ranging from blue (predominantly star-forming) to red (predominantly quenched). The underlying distribution is computed using a two-dimensional kernel density estimate (KDE) with adaptive smoothing. The quenched fraction is calculated as the ratio of the KDE-weighted density of quenched galaxies to the total KDE-weighted density in each cell. Black contours indicate the underlying density of individual galaxies. The halo masses utilised in this figure are estimated using the calibrated relation between the sum of the stellar masses of the three most massive group galaxies and the group halo mass. This figure highlights the dependence of quenching on both stellar and halo mass, with the quenched fraction increasing towards higher masses in both dimensions.
The second application demonstrates the utility of the SHMR-based halo mass estimator for investigating the interplay between galaxy quenching, stellar mass, and environment. To ensure compatibility with the SHMR calibration detailed in Section 3.2.3, we applied a systematic offset of +0.13 dex to all observational stellar mass estimates, bringing them into alignment with the ProSpect like stellar masses used in the Shark calibration. This adjustment is essential for accurate halo mass estimation using Equation 11 (see Section 3.2.3 for more details). Figure 6 presents the distribution of the quenched fraction in the stellar mass–halo mass plane for the mock datasets, while Figure 7 shows the corresponding distributions for the observational dataset. The quenched fraction is defined as the fraction of galaxies with $\log\,\mathrm{sSFR}<-11.0~[\mathrm{yr}^{-1}]$, following the criteria established in Van Kempen et al. (2024), with observational SFR upper limits classified as quenched if they meet the W3 detection limit for their given distance, otherwise conservatively classified as star-forming (see Van Kempen et al. (2024) for full details regarding upper-limit handling).
@@ 5 Summary and Conclusions
The accurate estimation of dark matter halo masses for galaxy groups is a fundamental requirement for advancing our understanding of galaxy evolution and for leveraging group catalogues as cosmological probes. In this work, we have developed, calibrated, and systematically evaluated two complementary halo mass estimators—one based on the virial theorem and the other on the summed stellar mass of the most massive group galaxies—using a suite of state-of-the-art semi-analytic models (SAMs) and observationally motivated mock catalogues. Our primary aim was to provide robust, minimally model-dependent, and observationally practical methods for inferring group-scale halo masses, with well-quantified uncertainties and clear guidance on their applicability.
We began by motivating the need for improved halo mass estimators in the context of upcoming spectroscopic surveys and the limitations of existing approaches (Section 1). We then described the construction of our simulation and observational datasets, including the use of Shark, SAGE, and GAEA SAMs, and the generation of realistic mock light-cones (Section 2). The core of the paper focused on the calibration and validation of the two halo mass estimators: the calibrated virial theorem (Section 3.1) and the stellar-to-halo mass relation (SHMR; Section 3.2). Each method was rigorously tested across multiple SAMs, with performance quantified in terms of accuracy, uncertainty, and model dependence (Section 3.1.3, Section 3.2.3). Finally, we demonstrated the utility of these estimators in two key observational applications: the construction of the empirical halo mass function and the mapping of quenched fractions in the stellar mass–halo mass plane (Section 4).
@@
    Calibrated Virial Theorem: We present a robust, physically motivated calibration of the virial theorem for group-scale halo mass estimation, incorporating corrections for velocity dispersion and projected radius (Section 3.1). The calibrated estimator achieves negligible systematic bias (mean $\Delta\sim 0.01$ dex) and moderate scatter ($\sigma\sim 0.23$$0.25$ dex) across all tested SAMs, with minimal dependence on baryonic physics (Section 3.1.3, Figure 3).
@@
    Stellar-to-Halo Mass Relation (SHMR): We calibrate an empirical relation between the sum of the stellar masses of the three most massive group galaxies and the halo mass, optimised via Bayesian MCMC (Section 3.2). This method yields the highest precision among those tested, with typical scatter as low as $0.12$ dex in the primary calibration sample, but exhibits greater model dependence due to its sensitivity to the underlying baryonic physics and stellar mass assembly (Section 3.2.3, Figure 4).
@@
    Comparison Across Models: Both estimators were validated across Shark, SAGE, and GAEA, demonstrating robust performance. The virial theorem calibration is largely insensitive to the details of the SAM, while the SHMR shows greater variance, particularly in models with differing feedback and mass resolution (notably GAEA at low masses; Section 3.1.3, Section 3.2.3).
@@
    Observational Applications: We showcase two key applications: (i) the construction of the empirical halo mass function (HMF) using the calibrated virial theorem, demonstrating the feasibility of this approach for future spectroscopic surveys (Section 4, Figure 5); and (ii) the mapping of the quenched fraction in the stellar mass–halo mass plane using SHMR-based halo masses, revealing the joint dependence of quenching on both stellar and halo mass (Section 4, Figure 7).
@@
    Guidance for Future Work: For cosmological applications, such as HMF construction and parameter inference, the calibrated virial theorem is recommended due to its minimal model dependence. For studies focused on group-scale halo mass estimation and environmental effects, the SHMR provides the highest precision, provided reliable stellar mass measurements are available (Section 3.1.3, Section 3.2.3, Section 4).
@@ 6 Acknowledgements
M.E.C. is a recipient of an Australian Research Council Future Fellowship (project No. FT170100273) funded by the Australian Government. D.J.C. is a recipient of an Australian Research Council Future Fellowship (project No. FT220100841) funded by the Australian Government. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. GAMA is a joint European-Australasian project based around a spectroscopic campaign using the Anglo-Australian Telescope. The GAMA input catalogue is based on data taken from the Sloan Digital Sky Survey and the UKIRT Infrared Deep Sky Survey. Complementary imaging of the GAMA regions is being obtained by a number of independent survey programmes including GALEX MIS, VST KiDS, VISTA VIKING, WISE, Herschel-ATLAS, GMRT and ASKAP providing UV to radio coverage. GAMA is funded by the STFC (UK), the ARC (Australia), the AAO, and the participating institutions. The GAMA website is https://www.gama-survey.org/. Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 177.A-3016. This research has made use of python (https://www.python.org) and python packages: astropy (Astropy Collaboration et al., 2013, 2018, 2022), cmasher (van der Velden, 2020), emcee (Foreman-Mackey et al., 2013), matplotlib http://matplotlib.org/ (Hunter, 2007), NumPy http://www.numpy.org/ (van der Walt et al., 2011), Pandas (McKinney, 2010), and SciPy https://www.scipy.org/ (Virtanen et al., 2020).
@@ Appendix A Virial Theorem Fitting Parameters
The calibration of the virial theorem-based halo mass estimator was performed using a Bayesian Markov Chain Monte Carlo (MCMC) approach, enabling robust exploration of the posterior probability distributions for all model parameters. The likelihood function was constructed to minimise the scatter in $\log M_{\mathrm{halo}}$ at fixed predicted mass, and is defined as
@@
The resulting posterior distributions and covariances for the six free parameters are shown in Figure A8. The median values and 68% credible intervals for each parameter are:
$\alpha$ $=1.030^{+0.092}_{-0.079},$
$\sigma_{\mathrm{lim}}$ $=244.634^{+5.398}_{-5.348}~\mathrm{km~s}^{-1},$
$\beta$ $=0.213^{+0.122}_{-0.085},$
$R_{\mathrm{lim}}$ $=0.369^{+0.083}_{-0.055}~\mathrm{Mpc},$
$n_{1}$ $=-1.989^{+0.032}_{-0.032},$
$n_{2}$ $=-1.591^{+0.123}_{-0.124}.$
The corner plot in Figure A8 illustrates that all parameters are well-constrained, with physically plausible correlations and no significant degeneracies. This calibration provides a statistically robust foundation for the application of the virial theorem-based halo mass estimator to both simulated and observational group catalogues. For further details on the calibration methodology and performance, see Sections 3.1 and 3.1.3.

### v3

@@ 5 Observational Application of Halo Masses
The calibrated halo mass estimators developed in this work enable a range of new applications in the analysis of spectroscopic group catalogues. Here, we present two representative examples using the SGP group sample from VanKempen2024: (i) the construction of an empirical HMF using the MVT, and (ii) the mapping of the quenched fraction of galaxies in the stellar mass–halo mass plane using the sSHMR.
@@ 5.1 Halo Mass Function from Observational Groups
The HMF is a fundamental cosmological observable, that encodes information about the growth of structure and the underlying cosmological parameters (e.g., Press1974; Jenkins2001; Murray2013). In Figure 6, we present the empirical HMF for galaxy groups with $N_{\mathrm{m}}\geq 4$ members. The sample is divided into three subsamples: the SGP-2dF (red triangles) which uses 2dFGRS photometry across the entire sample, the G23-2dF region (green crosses) which uses 2dFGRS photometry in the G23 region, and the G23-GAMA (orange diamonds) which uses the GAMA photometry.
Figure 6: Empirical HMF for galaxy groups with three or more members, constructed from the group catalogue of VanKempen2024. The HMF is shown separately for groups in the SGP region dominated by 2dF coverage (red triangles), the G23 region with GAMA spectroscopy (orange diamonds), and the G23 region with 2dF spectroscopy (green crosses). Halo masses are estimated using the traditional and calibrated virial theorem method ($\log M_{\mathrm{halo,\,VT}}$ or $\log M_{\mathrm{halo,\,MVT}}$). Number densities are not corrected for survey volume or selection effects, due to the heterogeneous nature of the group sample and the challenges in defining a complete selection function. The solid black curve represents the analytic HMF prediction for the adopted cosmology of Shark, while the grey dashed line shows the empirical fit from Driver2022 based on GAMA5, SDSS5, and REFLEX II data. Error bars reflect Poisson uncertainties.
@@
Unlike the SMF, which is based on uniform measurements of individual galaxies, constructing the HMF requires consistent data for all members within each group and a well-defined selection function to enable $1/V_{\mathrm{max}}$ corrections. However, the observational group sample is heterogeneous, with the formation and detection of groups often drawing from multiple spectroscopic surveys that have varying completeness and magnitude limits (see VanKempen2024). This diversity makes it challenging to accurately determine the completeness for each group, and thus we do not apply the $1/V_{\mathrm{max}}$ correction to the observational HMF. The absence of this correction means that the HMF is not fully corrected for incompleteness, particularly at the low-mass end. Low-mass groups are more likely to be missed because their member galaxies often fall below the survey’s limiting magnitude, and even when detected, such groups are typically found only at very low redshifts, where the survey volume is small and statistical uncertainties are larger. As a result, the observed HMF underestimates the true abundance of low-mass haloes, leading to an artificial drop-off in this range. Therefore, the HMF presented here should be regarded as a demonstration of the methodology rather than a definitive measurement.
@@
To ensure reliable binning of the HMF, the provided HMFs includes groups with $N_{\mathrm{m}}\geq 3$ detected in each survey (SGP-2dF, G23-2dF, and G23-GAMA). For the SGP-2dF sample, an additional restriction to $z<0.08$ is imposed to minimise the effects of declining completeness at higher redshifts. The HMFs shown in Figure 6 are derived using both the standard (left panel: Equation 3) and MVT (right panel: Equation 6) estimators. As discussed in Section 4.1.1 and illustrated in Figure 4, the standard virial theorem estimator predicts an excess of very low-mass group haloes. When considering the observational survey volume and the presence of a local void at $z\sim 0.02$ (Driver2022b; VanKempen2024), such low-mass systems are not expected to be sampled within this dataset. This underestimation of halo masses by the virial theorem could be misinterpreted as a genuine feature in the absence of this prior knowledge. In contrast, the MVT corrects for this bias, yielding a halo mass function that is more consistent with theoretical and practical expectations, providing a more reliable representation of the underlying group halo mass distribution.
@@
The resulting HMFs are compared both to the analytic prediction for the adopted cosmology of Shark and to the empirical fit derived by Driver2022. The empirical fit from Driver2022 is based on halo masses drawn from three low-redshift group/cluster catalogues: the GAMA group catalogue of Robotham2011, the SDSS Data group catalogue of Tempel2017, and ROSAT-ESO Flux Limited X-ray Galaxy Cluster Survey (REFLEX II; Boehringer2013; Boehringer2014; Boehringer2017). Together these span the mass range from galaxy groups to massive clusters, enabling a continuous measurement of the $z=0$ HMF. It is important to note that the dataset used by Driver2022 to establish their empirical relation is substantially larger and homogenised compared the sample analysed in this work. Nevertheless, our non-idealised and considerably smaller dataset yields a HMF that aligns well with theoretical expectations at the high to moderate halo mass range ($\log M_{halo}>12.75~[M_{\odot}]$) and underscores the significant progress made in the development of accurate halo mass estimators. However, at lower halo masses, the observed HMFs exhibit clear signs of incompleteness, further emphasising the necessity for larger and more homogeneous samples to robustly characterise this regime and improve observational constraints on the low-mass end.
@@
Looking ahead, upcoming wide-area spectroscopic surveys such as WAVES (Driver2019) and 4HS (ENTaylor23) will provide the depth, completeness, and sky coverage required to construct high-precision HMFs over a wide mass range. With sufficiently large samples and accurate halo mass estimates, such surveys will enable the use of the HMF as a cosmological probe, constraining parameters such as $\Omega_{m}$, $\sigma_{8}$, and potentially the nature of dark energy (e.g., Murray2013; Bocquet2019). The principal limiting factor for future HMF-based cosmological analyses is likely to be the uncertainty in group identification and membership assignment, particularly at low multiplicity, rather than the precision of the halo mass estimator itself.
@@ 5.2 Quenched Fraction in the Stellar Mass–Halo Mass Plane
Figure 8: Distribution of quenched fraction in the stellar mass–halo mass plane for galaxy groups in the observational dataset of VanKempen2024. The colour scale represents the fraction of quenched galaxies (defined as those with $\log\,\mathrm{sSFR}<-11.0$) within each region of parameter space, ranging from blue (predominantly star-forming) to red (predominantly quenched). The underlying distribution is computed using a two-dimensional kernel density estimate (KDE) with adaptive smoothing. The quenched fraction is calculated as the ratio of the KDE-weighted density of quenched galaxies to the total KDE-weighted density in each cell. Black contours indicate the underlying density of individual galaxies. The halo masses used in this figure are estimated using the calibrated sSHMR between the sum of the stellar masses of the three most massive group galaxies and the group halo mass. This figure highlights the dependence of quenching on both stellar and halo mass, with the quenched fraction increasing towards higher masses in both dimensions.
The second application demonstrates the utility of the sSHMR estimator for investigating the interplay between galaxy quenching, stellar mass, and environment. To ensure compatibility with the sSHMR calibration detailed in Section 4.2.3, we applied a systematic offset of +0.13 dex to all observational stellar mass estimates, bringing them into alignment with the ProSpect like stellar masses used in the Shark calibration. This adjustment is essential for accurate halo mass estimation using Equation 11 (see Section 4.2.3 for more details). Figure 7 presents the distribution of the quenched fraction in the stellar mass–halo mass plane for the mock datasets, while Figure 8 shows the corresponding distributions for the observational dataset. The quenched fraction is defined as the fraction of galaxies with $\log\,\mathrm{sSFR}<-11.0~[\mathrm{yr}^{-1}]$, following the criteria established in VanKempen2024, with observational SFR upper limits classified as quenched if they meet the W3 detection limit for their given distance, otherwise conservatively classified as star-forming (see VanKempen2024 for full details regarding upper-limit handling).
@@ 6 Summary and Conclusions
Accurate estimation of dark-matter halo masses for galaxy groups is fundamental both for studies of galaxy evolution and for exploiting group catalogues as cosmological probes. In this work we developed, calibrated, and validated two complementary, observational halo-mass estimators: a modified virial-theorem estimator (MVT) and a summed stellar–halo mass relation (sSHMR) that uses the sum of the stellar masses of the three most massive group members, as the predictor. The estimators were constructed using realistic mock light cones and survey selections (Sections 4.1 & 4.2); both were calibrated on the fiducial Shark SAM and subsequently validated on SAGE and GAEA catalogues to quantify model dependence. Performance was quantified in terms of accuracy, uncertainty, and variations between models (Sections 4.1.3, 4.2.3); the MVT was designed to minimise model dependence while the sSHMR provides the lowest scatter where reliable stellar masses are available. Finally, the practical utility of the calibrated estimators was demonstrated in two observational applications, the empirical halo mass function and the mapping of quenched fractions in the stellar mass–halo mass plane (Section 5).
@@
    Modified Virial Theorem: We present a robust, physically motivated modification of the virial theorem for group-scale halo mass estimation, incorporating corrections for velocity dispersion and projected radius (Section 4.1). The calibrated estimator achieves negligible systematic bias (mean $\Delta\sim 0.01$ dex) and moderate scatter ($\sigma\sim 0.23$$0.25$ dex) across all tested SAMs, with minimal dependence on baryonic physics (Section 4.1.3, Figure 4).
@@
    summed Stellar-to-Halo Mass Relation: We calibrate an empirical relation between the sum of the stellar masses of the three most massive group galaxies and the halo mass, optimised via Bayesian MCMC (Section 4.2). This method yields the highest precision among those tested, with typical scatter as low as $0.12$ dex in the primary calibration sample, but exhibits greater model dependence due to its sensitivity to the underlying baryonic physics and stellar mass assembly (Section 4.2.3, Figure 5).
@@
    Comparison Across Models: Both estimators were callibrated to Shark and cross-validated with SAGE, and GAEA, demonstrating robust performance. The virial theorem calibration is largely insensitive to the details of the SAM, while the sSHMR shows greater variance, particularly in models with differing feedback and mass resolution (notably GAEA at low masses; Section 4.1.3, Section 4.2.3).
@@
    Observational Applications: We showcase two key applications: (i) the construction of the empirical HMF using the calibrated virial theorem, demonstrating the feasibility of this approach for future spectroscopic surveys (Section 5, Figure 6); and (ii) the mapping of the quenched fraction in the stellar mass–halo mass plane using sSHMR-based halo masses, revealing the joint dependence of quenching on both stellar and halo mass (Section 5, Figure 8).
@@
    Guidance for Future Work: For cosmological applications, such as HMF construction and parameter inference, the calibrated virial theorem is recommended due to its minimal model dependence. For studies focused on group-scale halo mass estimation and environmental effects, the sSHMR provides the highest precision, provided reliable stellar mass measurements are available (Section 4.1.3, Section 4.2.3, Section 5).
@@ 7 Acknowledgements
We thank the anonymous referee for helpful comments and suggestions that have improved the content and clarity of this paper. M.E.C. is a recipient of an Australian Research Council Future Fellowship (project No. FT170100273) funded by the Australian Government. D.J.C. is a recipient of an Australian Research Council Future Fellowship (project No. FT220100841) funded by the Australian Government. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. GAMA is a joint European-Australasian project based around a spectroscopic campaign using the Anglo-Australian Telescope. The GAMA input catalogue is based on data taken from the Sloan Digital Sky Survey and the UKIRT Infrared Deep Sky Survey. Complementary imaging of the GAMA regions is being obtained by a number of independent survey programmes including GALEX MIS, VST KiDS, VISTA VIKING, WISE, Herschel-ATLAS, GMRT and ASKAP providing UV to radio coverage. GAMA is funded by the STFC (UK), the ARC (Australia), the AAO, and the participating institutions. The GAMA website is https://www.gama-survey.org/. Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 177.A-3016. This research has made use of python (https://www.python.org) and python packages: astropy (Astropy13; Astropy18; Astropy2022), cmasher (cmasher), emcee (emcee), matplotlib http://matplotlib.org/ (Hunter07), NumPy http://www.numpy.org/ (Walt11), Pandas (Pandas), and SciPy https://www.scipy.org/ (Virtanen20).
@@ Appendix A Virial Theorem Fitting Parameters
The calibration of the MVT was performed using a Bayesian MCMC approach, enabling robust exploration of the posterior probability distributions for all model parameters. The likelihood function was constructed to minimise the scatter in $\log M_{\mathrm{halo}}$ at fixed predicted mass, and is defined as
@@
The resulting posterior distributions and covariances for the six free parameters are shown in Figure A9. The median values and 68% confidence intervals for each of the coefficients are are shown in Table A3.
Table A3: Best-fitting of the coefficients for the modified virial theorem relation, including their 16th and 84th percentile scatter.
Coefficient $16^{th}$ Percentile Median $84^{th}$ Percentile
$\alpha$ 0.951 1.030 1.122
$\sigma_{lim}[\mathrm{km/s}]$ 239.286 244.634 250.032
$n_{1}$ -2.021 -1.989 -1.957
$\beta$ 0.128 0.213 0.335
$R_{lim}[\mathrm{Mpc}]$ 0.314 0.369 0.452
$n_{2}$ -1.715 -1.591 -1.468
The corner plot in Figure A9 illustrates that all parameters are well-constrained, with physically plausible correlations and no significant degeneracies. This calibration provides a statistically robust foundation for the application of the virial theorem-based halo mass estimator to both simulated and observational group catalogues. For further details on the calibration methodology and performance, see Sections 4.1 and 4.1.3.